# Shahriar Noroozizadeh

### v2

@@ Abstract
Then, by analyzing a connection to Node2Vec, we demonstrate how the geometry stems from a spectral bias that—in contrast to prevailing theories—indeed arises naturally despite the lack of various pressures. This analysis also points out to practitioners a visible headroom to make Transformer memory more strongly geometric. We hope the geometric view of parametric memory encourages revisiting the default intuitions that guide researchers in areas like knowledge acquisition, capacity, discovery, and unlearning.
@@ 1 Introduction
When succinct high-level patterns explain the data, deep sequence models produce corresponding high-level representations, as seen in natural language [40, 130, 112] and arithmetic tasks [100, 120, 117, 59]. When no such patterns exist (e.g., as in the capitals of countries), models are said to default to a brute-force lookup, known as associative memory [139, 20]. These two narratives have so far roughly guided our understanding of how neural networks fit sequential data. We highlight a third behavior missing in this narrative: even when memorizing such incompressible atomic co-occurrences, a deep sequence model can produce powerful representations. This implies a distinct form of parametric memory that we term as geometric memory—glimpses of which can be found in recent observations [74, 85, 189]. In opposition to the well-studied associative memory, a geometric memory has synthesized “global” information not explicit in the local co-occurrences, enabling new forms of reasoning. From this behavior, we extract aspects of neural embedding geometries that are hard to explain, raising fundamental questions about memorization in deep sequence models. To these questions, we offer some preliminary answers.
@@
Our discussion starts at a seemingly tangential point. We begin by consolidating a fragmented set of recent demonstrations [85, 174, 43, 48, 189, 74] that the Transformer exhibits some level of implicit in-weights reasoning, i.e., reasoning over knowledge from the weights without emitting an explicit chain of thought. We sharpen these results by crafting a scenario where this ability is unexpected, plays out vividly, and can be cleanly isolated and analyzed. Specifically, we study path-finding on path-star graphs, a (symbolic) implicit reasoning task. The task was adversarially designed [13] to cause failure of next-token trained deep sequence models—Transformer [172] and Mamba [60] models alike. Whereas in the original task, the model is given the graph in-context, here we make the model memorize the graph’s edges in its weights. Where before the model spectacularly failed to learn path-finding even on small graphs, in our in-weights task the model succeeds even on massive graphs.
@@
Figure 1: Associative vs. geometric memory of models trained on various graphs. There are two dramatically different ways to memorize a dataset of atomic facts. The common view is of associative memory: entities are embedded arbitrarily, and co-occurrences are stored in weight matrices. (left). §2.4: In practice, we find a geometric memory: the learned embeddings of a Transformer (middle) reflect global structure inferred from the local co-occurrences in training data. §4: When associative memory is explicitly prohibited (by removing intermediate layers), as in a Node2Vec model (right), a more elegant geometry materializes. This points to a clear headroom to improve the geometric nature of a Transformer’s memory. Details of the Transformer architecture used for this visualization are provided in §B.2.2. Similar geometries for Mamba and neural networks are presented in §C.3.
@@
The success in the in-weights path-star task, we argue, is hard to reconcile within the associative view of parametric memory, a convenient and highly effective abstraction of neural network memory, popular in literature [20, 26, 78, 177, 51, 202, 27, 124, 176]. In this abstraction (see Section˜2.3), each entity is embedded randomly/near-orthogonally via a “key” function ${\bf{\Phi}}(\cdot)$, while associations are stored in what is effectively a (rotated) adjacency matrix, ${\bf{W}}_{\tt assoc}$. To perform a “lookup”, one computes the quantity ${\bf{\Phi}}(v)^{T}{\bf{W}}_{\tt assoc}{\bf{\Phi}}(u)$ which is large if and only if entities $u$ and $v$ co-occur during training. With such a data structure, our path-finding task requires composing the lookup operation $\ell$ times (for path length $\ell$). Unless there is step-wise supervision for each composition, learning the $\ell$-fold composition should be a daunting needle-in-the-haystack task demanding $\exp(\ell)$ time. By the specific design of our task, every possible form of step-wise guidance is eliminated. Yet our model appears to find the needle.
@@
This apparent paradox begins to be resolved by a preliminary observation in [78, 85, 189]: that the node embeddings $\bf{\Phi}$} reflect some useful notion of distance; and a separate finding in [74, 153]: that (two-hop) reasoning can emerge from “factorized” parameterizations. We flesh out a unified understanding of these observations and their implications. First, the existence of structured node embeddings implies that models can store atomic facts in an altogether distinct paradigm of parametric memory that we term as geometric memory. This memory is in dramatic contrast to the associative one. In the simplest form of the geometric view (see Section˜2.4), the associations are no longer stored verbatim in a lookup table ${\bf{W}}_{\tt assoc}$, but are (low-rank) factorized into the embeddings as ${\bf{\Phi}}_{\tt geom}(u)^{T}{\bf{\Phi}}_{\tt geom}(v)$. Our key insight is that these embeddings are no longer arbitrary. Rather, they are carefully arranged such that they encode global relationships without explicit supervision to do so: even if entities $u$ and $v$ never appeared in the same context, the dot-product ${\bf{\Phi}}_{\tt geom}(u)^{T}{\bf{\Phi}}_{\tt geom}(v)$ captures the model’s own notion of multi-hop distance between the pair. This global geometry can open up new forms of reasoning, since what seemed to be a hard-to-learn $\ell$-fold composition of local associations is now approximated as an easy-to-learn spatial navigation task. We contrast the associative (Section˜2.3) and geometric (Section˜2.4) views of memory visually in Fig.˜1 and also in Table˜1.
@@
The observed geometry raises fundamental questions. Most importantly, there must be competition between the two parametric memories, both equally valid solutions to the training objective; why does the geometric prevail over the associative? To some readers, a geometric bias may seem familiar and intuitive at first sight, but we isolate aspects that cannot be easily explained: the global geometry is learned even when there is no “global" supervision (e.g., a path-finding objective), even when there is no rank constraint, even when the geometry is much harder to find, and even when it is just as succinct as a lookup. Thus, typical pressures from the supervision, architecture or optimization do not explain why a highly non-trivial geometry is synthesized. We term this the memorization puzzle. Towards understanding this, we reduce the setup to a minimal Node2Vec architecture, where we find that global geometries emerge from well-known spectral biases in such architectures. But, in contrast to prevailing theories (e.g., Levy and Goldberg [93]; see §5.5), we identify how the spectral bias arises naturally from cross-entropy loss minimization, independent of typically-assumed architectural or optimizational pressures. This gives us a preliminary insight into the natural rise of a geometric memory, albeit in a much simpler model, leaving open a foundational question for deep sequence models. Various practical questions open up too, which we discuss below.
@@ 1.1 Implications
Although the evidence of implicit reasoning is so far limited to symbolic tasks, we believe that the geometry we isolate is a clean nucleus of well-established geometries in language modeling [40, 112, 65] and arithmetic tasks [100, 120, 117, 59]. These insights suggest broader directions, both practical and theoretical. First, we find that the embeddings learned by more naive Node2Vec models are more strongly geometric than those of Transformers; in hindsight, this points to a well-specified headroom for making Transformer memory more geometric and less associative in practice. If the geometric bias can be improved in natural language tasks, it would also benefit natural implicit reasoning tasks where results have so far been mixed [135, 21, 185, 186, 16]. Since geometric memory encodes global relationships, it could help discover novel connections between information scattered in a large pretraining set [187] and thus, some types of creativity [22, 45, 119]. On the flip side, the interdependencies in geometric storage may impose limits on knowledge editing, unlearning and accurate retrieval; it may also hallucinate associations [74] (cf. “illusory correlations” in [153]). Another implication of our study is a support for why parametric memory may be superior to in-context memory, echoing Wang et al. [174], Geerts et al. [48]. Finally, the gap between the Transformer and Node2Vec geometries may also be of interest to practitioners in retrieval systems, when choosing between modern generative retrieval models [166, 179, 141] and traditional dual-encoder models [52, 73].
Many foundational directions follow. Foremost are the questions of how associative and geometric memory compete under gradient descent and what aspects of the data or optimization are ideal for the geometry to arise. The examples and analyses in this work may act as sandboxes to study a variety of geometric empirical phenomena: the emergence of “world models” [44, 65, 151], linear representations and superposition in interpretability [112, 130, 40] and why different models share similar representations, i.e., the Platonic representation hypothesis [92, 75]—and what these representations are. Finally, we speculate that the associative view forms the default unstated set of intuitions that guide research in numerous areas such as knowledge acquisition, discovery, unlearning, reasoning, storage capacity and scaling laws. The geometric view may inspire researchers to revisit, spell out and widen these latent intuitions. For instance, one may re-examine capacity [106, 81, 105, 80, 87] and scaling laws [10, 115, 144, 128] under a geometric lens and find very different answers.
@@ 1.2 Summary of contributions
    We isolate a clean instance of implicit in-weights reasoning in deep sequence models that succeeds even without any form of stepwise supervision. This isolates a behavior that is hard to explain within the predominant associative view of parametric memory. (§2)
@@
    We demonstrate why the geometric memory is surprising: the emergence of geometric over associative memory cannot be attributed to obvious architectural, optimizational or supervisory pressures. With this, we formulate a memorization puzzle in sequence modeling. (§3)
@@
    We connect the global geometry to the spectral bias of Node2Vec dynamics. We empirically intuit how it emerges without typically-assumed pressures, when minimizing the cross-entropy loss. This makes progress towards an open question in Node2Vec. We also highlight significant headroom in the embedding geometry of the current architectures. (§4)
@@
Simplest neural implementation (of the learned similarity between input token, node $u$ and next-token, node $v$, denoted by $f(u)[v]$) $f(u)[v]={\bf{\Phi}}(u)^{T}{\bf{W}}_{\tt assoc}{\bf{\Phi}}(v)$ where ${\bf{W}}_{\tt assoc}\approx\underset{(u,v)\in\mathcal{G}}{\sum}{\bf{\Phi}}(u)\otimes{\bf{\Phi}}(v)$} and ${\bf{\Phi}}(u)\cdot{\bf{\Phi}}(v)\approx 0$ for $u\neq v$. $f(u)[v]={\bf{\Phi}}_{\tt geom}(u)^{T}{\bf{\Phi}}_{\tt geom}(v)$ where ${\bf{\Phi}}_{\tt geom}$ corresponds to (top) eigenvectors of graph Laplacian.
Relationship to graph $\mathcal{G}$ ${\bf{W}}_{\tt assoc}$ can be thought of as adjacency matrix (up to a rotation) ${\bf{\Phi}}_{\tt geom}$ can be thought of as low-rank factorization of adjacency matrix; alternatively, the eigenvectors in ${\bf{\Phi}}_{\tt geom}$ are also components dominant in multi-hop adjacency matrix, $\bf{A}$$^{\ell}$ for large $\ell$.
Global information Only local edge information is readily available i.e., $f(u)[v]$ high only for $(u,v)\in E$ Global information readily available i.e., $f(u)[v]$ is proportional to multi-hop distance between $u$ and $v$ in $\mathcal{G}$
@@
Description length Number of edges, $|E|$ Number of vertices times the embedding dimensionality, |V|m|{V}|m
@@ 2 Implicit in-weights reasoning is learned
We investigate planning on a path-star graph, a graph designed to be adversarial towards next-token learning [13]. The task has a clear notion of a chain-of-thought, and a well-understood mechanism of failure for learning in-context reasoning. Thus, we hope to repurpose this to cleanly analyze a different type of reasoning—in-weights reasoning—in a way that was not possible in earlier studies of in-weights reasoning.
Background. The path-star topology (Fig.˜2) from B&N’24 is a tree graph where multiple disjoint uniform-length paths branch outwards from the root. In other words, all internal nodes but the root have exactly one child. In the original version of this task [B&N’24], the model is given a fresh example of this graph in context as an adjacency list (with randomized node labeling and edge ordering, but fixed topology), along with a leaf node specified as goal. The model must spell out the nodes in the unique path from the root to the goal in that order. The solution is to notice a simple right-to-left structure: take the unique path back from the leaf node and reverse it. Indeed, this is the implicit chain-of-thought required before emitting the first token.
Yet, on this simple task, left-to-right next-token learners are known to fail in-distribution even on tiny graphs. This failure unfolds in two stages during optimization: (1) On all but the first token, the model fits the data trivially: simply predict the token as the unique child of the previous ground-truth token that is revealed in the context. Then, this starves the first, key decision-making step of gradients from the rest of the path. (2) Consequently, the first token must be learned in isolation, which becomes a computationally hard learning $\ell$-hop composition problem (as detailed later). At test-time, this model defaults to guessing a random first token and continuing along that wrong path. We reproduce this failure in Fig.˜5 and elaborate on it in §A.
@@ 2.1 The in-weights path-star task
Task definition. For our study, we consider a task inspired by the above task, yet fundamentally different. In our task, the graph is stored in the model weights rather than presented in context (Fig.˜3). Here, all examples are generated from a (very large) fixed graph $\mathcal{G}$ (of degree $d$ and path length $\ell$), rather than a (small) fresh graph per example. The model is made to memorize the full graph in its weights through “edge-memorization examples”, where the input is some node $v$, and the next-token target is an adjacent node $v^{\prime}$. Separately, path-finding examples are generated from the same graph by picking as input a random leaf node ${v}_{\tt{leaf}}$ (a single token) and as target, the unique root-goal path (a sequence from ${v}_{\tt{root}}$ to ${v}_{\tt{leaf}}$). The model is trained for path-finding on only a subset of such leaves from our fixed graph, and is tested on the remaining “unseen” leaves. By “unseen”, we mean that the paths to these leaves are never seen end-to-end; only the constituent edges are seen individually, which doesn’t trivialize the path-finding task. (Besides this path-star topology, we consider some harder variants, deferred to §C.2.)
Experimental setup. We use a from-scratch, decoder-only Transformer (GPT-mid) [138]; we corroborate all our findings on Mamba in §C.1. For the most stable results, we interleave the edge-memorization and path-finding examples during training as in Fig.˜3 and also use pause tokens [57]. We found that it is important to provide both forward and reverse edges for edge-memorization in order to dodge the reversal curse [136, 9], but this may not be necessary for smaller graphs; see §D.1. This reverse augmentation is not given for the path-finding examples. Details about the formatting and the hyperparameters are in §B, followed by additional analyses in §D.

### v3

@@ Abstract
Then, by analyzing a connection to Node2Vec, we demonstrate how the geometry stems from a spectral bias that—in contrast to prevailing theories—indeed arises naturally despite the lack of various pressures. This analysis also points out to practitioners a visible headroom to make Transformer memory more strongly geometric. We hope the geometric view of parametric memory encourages revisiting the default intuitions that guide researchers in areas like knowledge acquisition, capacity, discovery, and unlearning.  [^1] Code for reproduction is at:
https://github.com/shahriarnz14/geometric_memory.
Geometric Memory, ICML
@@ 1 Introduction
When succinct high-level patterns explain the data, deep sequence models produce corresponding high-level representations, as often witnessed in natural language [44; 140; 121] and arithmetic tasks [108; 129; 126; 64]. When no such patterns exist (e.g., as in the capitals of countries), models are said to default to a brute-force lookup, known as associative memory [150; 22]. These two narratives have so far roughly guided our understanding of how neural networks fit sequential data. We highlight a third behavior overlooked in this narrative: even when memorizing such incompressible atomic co-occurrences, a deep sequence model can produce powerful representations. This implies a distinct form of parametric memory that we term as geometric memory—glimpses of which can be found scattered in recent observations [91; 134; 80; 200]. In opposition to the well-studied associative memory, a geometric memory has synthesized “global” information not explicit in the local co-occurrences, enabling new forms of reasoning. From this behavior, we extract aspects of neural embedding geometries that are hard to explain, raising fundamental questions about memorization in deep sequence models. To these questions, we offer some preliminary answers.
@@
To understand parametric memory, we study simple tasks involving graph memorization. These are tasks where the model is made to memorize the edge bigrams of an underlying graph in the model’s weights. As a starting point, we consolidate a fragmented set of recent demonstrations [91; 184; 47; 53; 200; 80] that, over such graphs, the Transformer can do some level of implicit in-weights reasoning, i.e., reasoning over parametric knowledge without emitting an explicit chain of thought. We sharpen these results by crafting a scenario where this ability is unexpected, plays out vividly, and can be cleanly isolated and analyzed. Specifically, we study path-finding on path-star graphs, a (symbolic) implicit reasoning task. The task was adversarially designed [14] to cause failure of next-token trained deep sequence models—Transformer [182] and Mamba [66] models alike. Whereas in the original task, the model is given the graph in-context, here we make the model memorize the graph’s edges in its weights. Where before the model spectacularly failed to learn path-finding even on small graphs, in our in-weights task the model succeeds even on massive graphs.
@@
Figure 1: Associative vs. geometric memory of models trained on various graphs. There are two dramatically different ways to memorize a dataset of atomic facts. The common view is of associative memory: entities are embedded arbitrarily, and co-occurrences are stored in weight matrices. (left). §2.4: In practice, we find a geometric memory: the learned embeddings of a Transformer (middle) reflect global structure inferred from the local co-occurrences in training data. §4: When associative memory is explicitly prohibited (by removing intermediate layers), as in a 2-layer, Node2Vec-style model (right), a more elegant geometry materializes. This points to a clear headroom to improve the geometric nature of a Transformer’s memory. Details of the Transformer architecture used for this visualization are provided in §B.2.2. Similar geometries for Mamba and neural networks are presented in §C.3. (Note: While these visualizations can tell us if a geometry exists, they do not definitively affirm how global the geometry is; for that, we recommend other tests such as examining the inner products. We discuss this in §B.5.)
@@
The success in the in-weights path-star task, we argue, is hard to reconcile within the associative view of parametric memory, a convenient and highly effective abstraction of neural network memory, popular in literature [22; 28; 84; 187; 56; 215; 29; 133; 186]. In this abstraction (see Section˜2.3), each entity is embedded randomly/near-orthogonally via a “key” function ${\bf{\Phi}}(\cdot)$, while associations are stored in what is effectively a (rotated) adjacency matrix, ${\bf{W}}_{\tt assoc}$. To perform a “lookup”, one computes the quantity ${\bf{\Phi}}(v)^{T}{\bf{W}}_{\tt assoc}{\bf{\Phi}}(u)$ which is large if and only if entities $u$ and $v$ co-occur during training. With such a data structure, our path-finding task requires composing the lookup operation $\ell$ times (for path length $\ell$). Unless there is step-wise supervision for each composition, learning the $\ell$-fold composition should be a daunting needle-in-the-haystack task demanding $\exp(\ell)$ time. By the specific design of our task, every possible form of step-wise guidance is eliminated. Yet our model appears to find the needle.
@@
This apparent paradox begins to be resolved by a preliminary observation in [84; 91; 134; 200]: that the node embeddings $\bf{\Phi}$} reflect some useful notion of distance; and a separate finding in [80; 163]: that (two-hop) reasoning can emerge from “factorized” parameterizations. We flesh out a unified understanding of these observations and their implications. First, the existence of structured node embeddings implies that models can store atomic facts in an altogether distinct paradigm of parametric memory that we term as geometric memory. This memory is in dramatic contrast to the associative one. In the simplest form of the geometric view (see Section˜2.4), the associations are no longer stored verbatim in a lookup table ${\bf{W}}_{\tt assoc}$, but are (low-rank) factorized into the embeddings as ${\bf{\Phi}}_{\tt geom}(u)^{T}{\bf{\Phi}}_{\tt geom}(v)$. Our key insight is that these embeddings are no longer arbitrary. Rather, they are carefully arranged such that they encode global relationships without explicit supervision to do so: even if entities $u$ and $v$ never appeared in the same context, the dot-product ${\bf{\Phi}}_{\tt geom}(u)^{T}{\bf{\Phi}}_{\tt geom}(v)$ captures the model’s own notion of multi-hop distance between the pair. This global geometry can open up new forms of reasoning, since what seemed to be a hard-to-learn $\ell$-fold composition of local associations is now approximated as an easy-to-learn spatial navigation task. We contrast the associative (Section˜2.3) and geometric (Section˜2.4) views of memory visually in Fig.˜1 and also in Table˜1.
@@
The observed geometry raises fundamental questions. Most importantly, there must be competition between the two parametric memories, both equally valid solutions to the training objective; why does the geometric prevail over the associative? To some readers, a geometric bias may seem familiar and intuitive at first sight, but we isolate aspects that cannot be easily explained: the global geometry is learned even when there is no “global" supervision (e.g., a path-finding objective), even when there is no rank constraint, even when the geometry is much harder to find, and even when it is just as succinct as a lookup. Thus, typical pressures from the supervision, architecture or optimization do not explain why a highly non-trivial geometry is synthesized. We term this the memorization puzzle. Towards understanding this, we reduce the setup to a minimal two-layer, Node2Vec-style architecture [65], where we find that global geometries emerge from well-known spectral biases in such architectures. But, in contrast to prevailing theories (e.g., Levy and Goldberg [100]; see §5.5), we identify how the spectral bias arises naturally from cross-entropy loss minimization, independent of typically-assumed pressures. While this gives us a preliminary insight into the natural rise of a geometric memory in a simple model, we leave open a foundational question for deep sequence models.
@@ 1.1 Implications
Although the evidence of implicit reasoning is so far limited to symbolic tasks, we believe that the geometry we isolate is a clean nucleus of well-established geometries in language modeling [44; 121; 71] and arithmetic tasks [108; 129; 126; 64]. These insights suggest broader directions, both practical and theoretical. First, we find that the embeddings learned by more naive Node2Vec models are more strongly geometric than those of Transformers; in hindsight, this points to a well-specified headroom for making Transformer memory more geometric and less associative in practice. If the geometric bias can be improved in natural language tasks, it would also benefit natural implicit reasoning tasks where results have so far been mixed [146; 23; 196; 197; 17]. Since geometric memory encodes global relationships, it could help discover novel connections between information scattered in a large pretraining set [198] and thus, some types of creativity [24; 49; 128]. On the flip side, the interdependencies in geometric storage may impose limits on knowledge editing (e.g., causing effects such as representation shattering [134]), unlearning and accurate retrieval; it may also hallucinate associations [80] (cf. “illusory correlations” in [163]). Another implication of our study is a support for why parametric memory may be superior to in-context memory, echoing Wang et al. [184]; Geerts et al. [53]. Finally, the gap between the Transformer and Node2Vec geometries may also be of interest to practitioners in retrieval systems, when choosing between modern generative retrieval models [176; 189; 152] and traditional dual-encoder models [57; 79].
Many foundational directions follow. Foremost are the questions of how associative and geometric memory compete under gradient descent and what aspects of the data or optimization are ideal for the geometry to arise. The examples and analyses in this work act as minimal and tractable sandboxes (with closed form solutions) to study a variety of complex geometric empirical phenomena: grokking [145], the emergence of “world models” [48; 71; 161], linear representations and superposition in interpretability [121; 140; 44] and why different models share similar representations, i.e., the Platonic representation hypothesis [99; 81]—and what these representations are. Finally, we speculate that the associative view forms the default unstated set of intuitions that guide research in numerous areas such as knowledge acquisition, discovery, unlearning, reasoning, storage capacity and scaling laws. The geometric view may inspire researchers to revisit, spell out and widen these latent intuitions. For instance, one may re-examine capacity [115; 87; 114; 86; 93] and scaling laws [10; 124; 155; 138] under a geometric lens and find very different answers.
@@ 1.2 Summary of contributions
    We devise a clean instance of implicit in-weights reasoning in deep sequence models that succeeds even without any form of step-wise supervision. This isolates a behavior that is hard to explain within the predominant associative view of parametric memory. (§2)
@@
    We demonstrate why the geometry is surprising: the emergence of geometric over associative memory cannot be attributed to obvious architectural, optimizational or supervisory pressures. With this, we formulate a memorization puzzle in sequence modeling. (§3)
@@
    We connect the global geometry to the spectral bias of 2-layer Node2Vec models. We empirically intuit how it emerges without typically-assumed pressures, when minimizing the cross-entropy loss. This makes progress towards an open question in Node2Vec. We also highlight significant headroom in the embedding geometry of the current architectures. (§4)
@@
Simplest neural implementation (of the learned similarity between input token, node $u$ and next-token, node $v$, denoted by $f(u)[v]$) $f(u)[v]={\bf{\Phi}}(u)^{T}{\bf{W}}_{\tt assoc}{\bf{\Phi}}(v)$ where ${\bf{W}}_{\tt assoc}\approx\underset{(u,v)\in\mathcal{G}}{\sum}{\bf{\Phi}}(u)\otimes{\bf{\Phi}}(v)$} and ${\bf{\Phi}}(u)\cdot{\bf{\Phi}}(v)\approx 0$ for $u\neq v$. $f(u)[v]={\bf{\Phi}}_{\tt geom}(u)^{T}{\bf{\Phi}}_{\tt geom}(v)$ where ${\bf{\Phi}}_{\tt geom}$ corresponds to eigenvectors of graph Laplacian (potentially, only the top eigenvectors).
Relationship to graph $\mathcal{G}$ ${\bf{W}}_{\tt assoc}$ can be thought of as adjacency matrix (up to a rotation) ${\bf{\Phi}}_{\tt geom}$ can be thought of as a factorization of adjacency matrix, typically low-rank or skewed towards the top eigenvectors.
Global information Only local edge information is readily available (or via a probe). That is, the logit $f(u)[v]$ high only for $(u,v)\in E$ Global information readily available. Concretely, $f(u)[v]$ is proportional to multi-hop distance between $u$ and $v$ in $\mathcal{G}$ (but if factorization is full rank and not skewed towards the top eigenvectors, global information can still be accessed through a probe that focuses on the top directions). Intuitively, the top eigenvectors contain global information as they are also the components dominant in multi-hop adjacency matrix, $\bf{A}$$^{\ell}$ for large $\ell$.
@@
Description length Number of edges, $|E|$ (upto logarithmic factors) Number of vertices times the embedding dimensionality, |V|m|{V}|m
Contents
@@ 2 Implicit in-weights reasoning is learned
We investigate planning on a path-star graph (Fig.˜2), a graph designed to be adversarial towards next-token learning [B&N’24]. The task has a clear notion of a chain-of-thought, and a well-understood mechanism of failure for learning in-context reasoning. Thus, we repurpose this to cleanly analyze a different type of reasoning—in-weights reasoning—in a way that was not possible in earlier studies of in-weights reasoning.
Figure 2: Our in-weights path-star task. All train/test examples are generated from a fixed path-star graph: (i) edge memorization examples (covering all edges during training) and (ii) path-finding examples, partitioned into a non-overlapping train and test set. Fuller version in Fig.˜10.
@@ 2.1 The in-weights path-star task
Our task definition. The path-star topology is a tree graph where multiple disjoint uniform-length paths branch out from the root. For our task, all examples are generated from a fixed (giant) graph $\mathcal{G}$ of degree $d$ and path length $\ell$. Part of the training data consists of “edge-memorization” examples where the input prompt is some node $v$ and the target is an adjacent node $v^{\prime}$ in $\mathcal{G}$. These training examples cover all the edges so that the model memorizes all of $\mathcal{G}$ in its weights. Separately, we generate path-finding examples by picking as input prompt a random leaf node ${v}_{\tt{leaf}}$ and as target, the unique root-goal path (a sequence of tokens from ${v}_{\tt{root}}$ to ${v}_{\tt{leaf}}$). We train for path-finding on only a subset of such leaves, and test on the remaining “unseen” leaves, leaves whose root-to-leaf paths are never seen end-to-end—only the constituent edges are seen individually, which doesn’t trivialize the path-finding task. (We extend our experiments to harder graphs in §C.2.)
Our experimental setup. We make some notes about our setup. First, for the most stable results, we interleave the edge-memorization and path-finding examples during training as in Fig.˜10 and also use pause tokens [62]. We found that it is important to provide both forward and reverse edges for edge-memorization in order to dodge the reversal curse [147; 9], but this may not be necessary for smaller graphs; see §C.3.1. Reverse augmentation is not given for the path-finding examples. We use from-scratch, decoder-only Transformer (GPT-mid) [149] and corroborate all findings on Mamba in §C.1. The formatting and the hyperparameters are in §B, followed by additional analyses in §D.
Comparison to the original in-context task. Our task is inspired by, yet fundamentally different from the original task in B&N’24. There, the model was given in context, a fresh example of a path-star graph as an adjacency list (with randomized node labeling and edge ordering), along with a leaf node specified as goal (see Fig.˜11 for this original version). We however generate all examples from a fixed in-weights graph, and give no adjacency list in-context as we hope to understand how well the model recalls and reasons over information stored in the weights.
The adversarial design of the graph. The solution to the path-star task is to notice a simple right-to-left structure: take the unique path back from the leaf node and reverse it. Indeed, this is the implicit chain-of-thought required before emitting the first token. Despite this simplicity, in the in-context version of this task, left-to-right next-token learners are known to fail in-distribution even on tiny graphs. This unfolds in two stages during optimization: (1) On all but the first token, the model fits the data trivially: predict the token as the unique child of the previous ground-truth token that is revealed in the context. This deprives the first, key decision-making step of gradients from the rest of the path. (2) Consequently, the first token must be learned in isolation, which is a computationally hard $\ell$-hop composition-learning problem (as detailed later). At test-time, this model defaults to guessing a random first token and continuing along that wrong path. For reference, we elaborate on this in-context task in §A and reproduce this failure in Fig.˜13.
Figure 3: Success of Transformer in in-weights path-star task. (§2) On large path-star graphs $\mathcal{G}_{d,\ell}$, a next-token-trained Transformer achieves perfect or highly non-trivial accuracy, even when trained only on the first-token in isolation. the $Y$-axis is in $\log$ scale to compare against chance accuracy ($1/d$; dashed line). More detailed plots in Fig.˜12; analogous plots for Mamba in Fig.˜14. We invite the reader to contrast these in-weights task results with the in-context task ones in Fig.˜13.